OpenStudy (anonymous):

last one!! given the function: g(x)=4^(1-x) find g(-1)

jimthompson5910 (jim_thompson5910):

g(x)=4^(1-x) g(-1)=4^(1-(-1)) ... replace EVERY x with -1 g(-1)=4^(1+1) g(-1)=4^2 g(-1)=16

OpenStudy (anonymous):

put \[x=4^{1-y}\] solve for y

jimthompson5910 (jim_thompson5910):

oh you're looking for the inverse?

OpenStudy (anonymous):

@jim i am going to bet that this means \[g^{-1}(x)\]

myininaya (myininaya):

satellite likes to do his own thing

OpenStudy (anonymous):

hell if i know! lol, this stuff isn't my cup of tea. it ust said simplify with integer or fraction

jimthompson5910 (jim_thompson5910):

hmm sounds like simple function evaluation to me

jimthompson5910 (jim_thompson5910):

can you draw it?

OpenStudy (anonymous):

g(-1)=_____?

OpenStudy (anonymous):

get \[\log_4(x)=1-y\] \[y=1-\log_4(x)\] \[g^{-1}(x)=1-\log_4(x)\] good luck with no more math !!!

OpenStudy (anonymous):

then again if it is \[g(-1)\] then ignore me and use jim's answer

OpenStudy (anonymous):

g(-1)=_____?

OpenStudy (anonymous):

LOL.

jimthompson5910 (jim_thompson5910):

or can you post a screenshot?

OpenStudy (anonymous):

g(-1)=_____?

OpenStudy (anonymous):

I THINK YOU'RE ANSWER WAS RIGHT, JIM....16. THANK YOU BOTH!!!